OpenStudy (anonymous):

Waves in a microwave have a frequence of 2.5 GHz, what wavelength do they have? The answer should be 0.12m. What am I doing wrong here? Here's what I get: Energy = (6,6*10^-34)*(2,5*10^9) = 1,65*10^-24 J Wavelength = ((6.6*10^-34)*(1.6*10^-19)) / 81,65*10^-24) = 6,4*10^-29

OpenStudy (anonymous):

Last line SHOULD read: Wavelength = ((6.6*10^-34)*(1.6*10^-19)) / (1,65*10^-24) = 6,4*10^-29

OpenStudy (anonymous):

Use, Velocity = Frequency * Wavelength Take Velocity = Velocity of Light = 3*10^8 m Given Frequency = 2.5 GHz = 2.5 *10^9 Hz => Wavelength = Velocity / Frequency